OpenStudy (anonymous):

Which poses more of a danger to human health, very long wavelength radiation or very short wavelength radiation? and explain why?

OpenStudy (anonymous):

high

OpenStudy (anonymous):

\[\Huge \color{red}{\ast~~\ast}~~\color{blue}{\ast~~\ast}~~\color{green}{\ast~~\ast}~~\color{orange}{\ast~~\ast}~~\color{purple}{\ast~~\ast}~~\color{salmon}{\ast~~\ast}~~\color{red}{\ast~~\ast}\] \[\LARGE \color{red}{\mathbb{WELCOME}}~~\color{orange}{\mathbb{TO}}~~\color{purple}{\mathbb{OPENSTUDY}}~~\color{blue}{!!}~~(:\] \[\Huge \color{red}{\ast~~\ast}~~\color{blue}{\ast~~\ast}~~\color{green}{\ast~~\ast}~~\color{orange}{\ast~~\ast}~~\color{purple}{\ast~~\ast}~~\color{salmon}{\ast~~\ast}~~\color{red}{\ast~~\ast}\]

OpenStudy (mos1635):

very long wavelength=> very low freqouency=> very low energy of photon => not harmful very short wavelength=> very high freqouency=> very high energy of photon => very harmful